# Near Optimal Locality Sensitive Orderings in Euclidean Space

### v2

@@ 2.1.1 From order on a grid, to orderings of the unit cube
Chan et al. [CHJ20] showed how to convert an order of a $t$-grid into an ordering of $[0,1)^{d}$. We now describe this construction in detail. Let $\mathcal{F}$ be the set of $\lambda$ $\varepsilon$-quadtrees of Lemma 2.3, and let $\mathcal{S}\subseteq[0,1)^{d}$ be the set of at most $d+1$ shifts of Lemma 2.4. Let $\mathfrak{O}$ be the set of orderings of the $t$-grid $\mathcal{G}(\mathcal{C},\mathcal{E})$, for some cube $\mathcal{C}$, as constructed by Lemma 2.5.
Next, consider an $\varepsilon$-quadtree $\mathcal{T}_{\varepsilon}\in\mathcal{F}$, a shift $\nu\in\mathcal{S}$, and an ordering $\sigma\in\mathfrak{O}$. This immediately induces an ordering over the points of $[0,1)^{d}$. Indeed, consider two points $p,q\in[0,1)^{d}$, and consider their shifted image $p^{\prime}=\nu+p$ and $q^{\prime}=\nu+q$. There is a unique least common ancestor (LCA) node $u\in\mathcal{T}_{\varepsilon}$ to the leafs where $p^{\prime}$ and $q^{\prime}$ would be stored in the quadtree. The node $u$ has $\mathcal{E}^{d}$ children, with $p^{\prime}$ and $q^{\prime}$ belonging to two different children, say $u_{p}$ and $u_{q}$. If $\sigma$ has $u_{p}$ before $u_{q}$, then we consider $p\prec_{\sigma}q$, and otherwise $p\succ_{\sigma}q$. Using standard bit operations, and the ordering $\sigma$ listed explicitly (which requires $O(\mathcal{E}^{d})$ space), this can be done in $O(\lambda)=O(\log\tfrac{1}{\varepsilon})$ time [CHJ20]. (The quadtree $\mathcal{T}_{\varepsilon}$ is used implicitly, so it is not computed explicitly.) Overall, the number of orderings this yield, is $\mathcal{O}(\left|{\mathfrak{O}}\right|\left|{\mathcal{S}}\right|\lambda)$. Given two points comparing their order according to a specific given order takes $O(\lambda)$ time. The resulting set of orderings of $[0,1)^{d}$ is of size $O(\mathcal{E}^{d}\log\mathcal{E})$. This implies the following theorem, see [CHJ20] for more details.
@@
There is a set $\Pi^{+}$ of $\mathcal{O}(\mathcal{E}^{d}\log\mathcal{E})$ orderings of $[0,1)^{d}$, such that for any two points $p,q\in[0,1)^{d}$ there is an ordering $\sigma\in\Pi^{+}$ defined over $[0,1)^{d}$, such that for any point $u$, with $p\prec_{\sigma}u\prec_{\sigma}q$, we have that either $\left\|{pu}\right\|\leq\varepsilon\left\|{pq}\right\|$ or $\left\|{qu}\right\|\leq\varepsilon\left\|{pq}\right\|$.
@@ 2.2 Coverings and packings
Consider a set $P\subseteq\mathbb{R}^{d}$, and a parameter $r$. A set $N\subseteq P$ is an $r$-covering of $P$, if for all $p\in P$, there exists a point $p^{\prime}\in N$, such that $\left\|{pp^{\prime}}\right\|\leq r$. A point set $N$ is $r$-separated if for any two distinct points $p,q\in N$, we have $\left\|{pq}\right\|>r$.
A subset $N\subseteq P$ that is both $r$-separated and an $r$-covering of $P$ is an $r$-packing [^1] Confusingly $r$-packings are also known as $r$-nets, but as the $r$-net is an overloaded concept, we use the alternative, hopefully less-confusing, term. of $P$.
For a finite set $P$ in $\mathbb{R}^{d}$, an $r$-packing can be computed in linear time [HR15] (for constant $d$). We need such a packing of the unit sphere.
@@
The set $N^{\prime}$ has (asymptotically) the right size. Converting it into an $r$-packing can be done in linear time using the algorithm mentioned above [HR15].
@@ 3.1 Projections
Let $R>\tau>0$ be parameters. One can compute, a set $\mathcal{D}$ of directions, such that for any two points $p,q\in\mathbb{R}^{d}$, such that $\left\|{pq}\right\|\leq R$, we have that there is a direction $v\in\mathcal{D}$, such that $\mathsf{d}_{v}\left({p,q}\right)\leq\tau$. The time to compute $\mathcal{D}$ is $O(\left|{\mathcal{D}}\right|)=O\bigl{(}(R/\tau)^{d-1}\bigr{)}$.
@@
Let $\mathcal{D}$ be an $\xi$-packing of the unit sphere $\mathbb{S}$, for $\xi=\min(\tau/R,1/4)$, computed by Lemma 2.9. As for correctness, consider the line $\mathcalb{l}$ connecting $p$ to $q$. Let $u$ be the direction vector of $\mathcalb{l}$, and let $v\in\mathcal{D}$ be the closest direction to $u$. By construction, $\left\|{uv}\right\|\leq\xi$. Let $\mathcalb{g}$ be the line passing through $p$ in the direction of $v$, and let $q^{\prime}$ be the projection of $q$ to this line, see Figure 3.1.
Figure 3.1:
Clearly, $\mathsf{d}_{v}\left({p,q}\right)=\left\|{qq^{\prime}}\right\|=\left\|{pq}\right\|\sin\angle qpq^{\prime}\leq R\sin\xi\leq R\xi\leq R\frac{\tau}{R}=\tau,$ since $\left\|{pq}\right\|\leq R$, $\sin x\leq x$ for $x$ positive, $\xi>0$, and $\xi\leq\tau/R$.
@@ 3.2 Constructing the ordering
    $\left\|{\mathsf{c}\left({\square_{3}}\right)\mathsf{c}\left({\square_{4}}\right)}\right\|\leq\left\|{\mathsf{c}\left({\square_{1}}\right)\mathsf{c}\left({\square_{2}}\right)}\right\|$,
@@
Here $\mathsf{c}\left({\square}\right)$ denotes the center point of normal-$\square$.
@@
We now describe this in more detail, for a fixed $v$ and $i$. This fix a set $N_{v}^{i}$. For $j=1,\ldots,\left|{N_{v}^{i}}\right|$, let $p_{j}$ be a point of $N_{v}^{i}$ not handled yet. The algorithm computes the oriented line $\mathcalb{l}_{j}$ passing through $p_{j}$ in the direction of $v$. Let $Z_{j}$ be the set of all the cells of $\mathcal{G}$ that $\mathcalb{l}_{j}$ stabs, and their center is in distance at most $\tau$ from $\mathcalb{l}_{j}$. All such cells intersect $\mathcalb{l}$ in a “long” interval of length $\geq\zeta-2\tau\geq(3/4)\zeta\geq 6\tau$ (that is, all the cells of $Z_{j}$ have their centers “almost” on $\mathcalb{l}_{j}$). Thus, the distance between the centers of the cells of $Z_{j}$ in the original space, or the distance between their corresponding projections on $\mathcalb{l}_{j}$ are the same up to ±$\pm 2\tau$. In particular, we sort the cells of $Z_{j}$ according to the order of their projected centers along $\mathcalb{l}_{j}$, and append them (in this sorted order) to the computed order. This process is illustrated in Figure 3.2. Note, that in the end of this process, most grid cells of $\mathcal{G}$ are not included in the computed order. We append all these yet unordered cells in an arbitrary fashion to the end of the ordering. Let $\sigma(v,i)$ denote the resulting ordering of the cells of $\mathcal{G}$.
@@ 3.3 An ordering with a gap
For two sets $X,Y\subseteq\mathbb{R}^{d}$, let $\mathsf{d}\left({X,Y}\right)=\min_{x\in X,y\in Y}\left\|{xy}\right\|$.
@@
and $N=\left|{\mathcal{G}_{T}}\right|=\Theta\bigl{(}(t/\beta)^{d})=\Theta\bigl{(}(t/\alpha)^{d}).$ We partition every cell normal-$\square$ of $\mathcal{G}_{T}$ into a secondary grid $\mathcal{G}_{\square}=\mathcal{G}(\square,\beta)$, see Figure 3.3. After applying this to all the grid cells of $\mathcal{G}_{T}$, this results in the original grid $\mathcal{G}$. Let $\mathfrak{O}_{T}$ be the set of orderings, of Lemma 2.5, of $\mathcal{G}_{T}$ that has the property that every two cells of $\mathcal{G}_{T}$ are adjacent in one of the orderings, where $\left|{\mathfrak{O}_{T}}\right|=O(N)$. Let $\mathfrak{O}_{B}$ be the set of orderings for the original grid $\mathcal{G}$ from Lemma 3.4. Consider an order $\sigma_{T}\in\mathcal{G}_{T}$ and an order $\sigma_{B}\in\mathfrak{O}_{B}$. They induce an ordering of the cells of $\mathcal{G}$ as follows – the top ordering $\sigma_{T}$ orders the cells of $\mathcal{G}$ into secondary blocks, where each secondary block is all the cells of $\mathcal{G}$ that lie in a single cell of $\mathcal{G}_{T}$. Every such block is then sorted using the order $\sigma_{B}$. this defines a new set of orders $\mathfrak{O}$ over $\mathcal{G}$ of size $O(\left|{\mathfrak{O}_{T}}\right|\cdot\left|{\mathfrak{O}_{B}}\right|)=O(t^{2d-1}/\alpha^{d})$.
@@ 3.4 Result
Let $\varepsilon,\gamma\in(0,1/2]$ be fixed constants, such that $\gamma\geq\varepsilon$. Then, there is a set $\Pi$ of $m=O\bigl{(}(\mathcal{E}^{d-1}\log\mathcal{E})/\gamma^{d}\bigr{)}$ orderings of $[0,1)^{d}$, such that for any two points $p,q\in[0,1)^{d}$, there is an ordering $\sigma\in\Pi$ defined over $[0,1)^{d}$, such that
@@
    for all points $u$, with $p\prec_{\sigma}u\prec_{\sigma}q$, we have $\mathsf{d}\left({u,pq}\right)\leq\varepsilon\left\|{pq}\right\|$ and $\mathsf{d}\left({u,\{p,q\}}\right)\leq\gamma\left\|{pq}\right\|$,
@@
    for all points $u,v$ with $p\prec_{\sigma}u\prec_{\sigma}v\prec_{\sigma}q$, we have $\left\|{uv}\right\|\leq(1+\varepsilon)\left\|{pq}\right\|$.
@@
[…]
Let $\mathfrak{O}=\mathfrak{O}\left({\mathcal{E},\alpha,d}\right)$ be the set of orderings of $\mathcal{G}(\mathcal{C},\mathcal{E})$ as defined by Lemma 3.5. As descried in Section 2.1.1, the set $\mathfrak{O}$ induces a set $\Pi=\Pi(\varepsilon^{\prime},\alpha)$ of orderings of $[0,1)^{d}$, where
@@
So consider any two points $p,q\in[0,1)^{d}$. There is a shift $\nu\in\mathcal{S}$, and an $\varepsilon^{\prime}$-quadtree $\mathcal{T}_{\varepsilon}\in\mathcal{F}$, such that the LCA of $p$ and $q$ in $\mathcal{T}_{\varepsilon}$ is a node $w$, such that $\mathrm{sidelen}\left({\square_{w}}\right)\leq(d+1)\left\|{pq}\right\|$, see Lemma 2.4, where $\mathrm{sidelen}\left({\square}\right)=\mathrm{diam}\left({\square}\right)/\sqrt{d}$ is the sidelength of normal-$\square$. In particular, let $\zeta=\mathrm{sidelen}\left({\square_{w}}\right)/\mathcal{E}$ be the sidelength of the subcells of the grid of $\mathcal{T}_{\varepsilon}$ for $\square_{w}$. Observe that $\left\|{pq}\right\|\geq\mathcal{E}\zeta/(d+1)$. Let $\square_{1}$ and $\square_{2}$ be the cells of the two children of $w$ that contains $p$ and $q$, respectively. In particular, we have that
@@
Computing the orderings of Theorem 3.7 can be done in $\mathcal{E}^{O(d)}$ time, and since this is a preprocessing stage, and conceptually we consider $\varepsilon$ to be a constant, we ignore this preprocessing time in our statement. This results in $m=O\bigl{(}(\mathcal{E}^{d-1}\log\mathcal{E})/\gamma^{d}\bigr{)}$ orderings, each order requires $O(\mathcal{E}^{d})$ space to store it. In the following, $\gamma$ is a fixed “large” constant (e.g., $\gamma=1/8$), so the overall space to store these orderings is $O(\mathcal{E}^{2d-1}\log\mathcal{E})$. Given an order $\sigma$ and two points, comparing the two points according to $\sigma$ requires answering a single LCA query on a quadtree, see Section 2.1.1. With the appropriate bit operations such a query can be carried out in $O(\log\mathcal{E})$ time [CHJ20].

### v3

@@ 2.2.1 From order on a grid, to orderings of the unit cube
Chan et al. chj-lota-20 showed how to convert an order of a $t$-grid into an ordering of $[0,1)^{d}$. We now describe this construction in detail. Let $\mathcal{F}$ be the set of $\lambda$ $\varepsilon$-quadtrees of Lemma 2.4, and let $\mathcal{S}\subseteq[0,1)^{d}$ be the set of at most $d+1$ shifts of Lemma 2.5. Let $\mathfrak{O}$ be the set of orderings of the $t$-grid $\mathcal{G}(\mathcal{C},\mathcal{E})$, for some cube $\mathcal{C}$, as constructed by Lemma 2.6.
Next, consider an $\varepsilon$-quadtree $\mathcal{T}_{\varepsilon}\in\mathcal{F}$, a shift $\nu\in\mathcal{S}$, and an ordering $\sigma\in\mathfrak{O}$. This immediately induces an ordering over the points of $[0,1)^{d}$. Indeed, consider two points $p,q\in[0,1)^{d}$, and consider their shifted image $p^{\prime}=\nu+p$ and $q^{\prime}=\nu+q$. There is a unique least common ancestor (LCA) node $u\in\mathcal{T}_{\varepsilon}$ to the leafs where $p^{\prime}$ and $q^{\prime}$ would be stored in the quadtree. The node $u$ has $\mathcal{E}^{d}$ children, with $p^{\prime}$ and $q^{\prime}$ belonging to two different children, say $u_{p}$ and $u_{q}$. If $\sigma$ has $u_{p}$ before $u_{q}$, then we consider $p\prec_{\sigma}q$, and otherwise $p\succ_{\sigma}q$. Using standard bit operations, and the ordering $\sigma$ listed explicitly (which requires $O(\mathcal{E}^{d})$ space), this can be done in $O(\lambda)=O(\log\tfrac{1}{\varepsilon})$ time chj-lota-20. (The quadtree $\mathcal{T}_{\varepsilon}$ is used implicitly, so it is not computed explicitly.) Overall, the number of orderings this yield, is $\mathcal{O}(\left|{\mathfrak{O}}\right|\left|{\mathcal{S}}\right|\lambda)$. Given two points comparing their order according to a specific given order takes $O(\lambda)$ time. The resulting set of orderings of $[0,1)^{d}$ is of size $O(\mathcal{E}^{d}\log\mathcal{E})$. This implies the following theorem, see chj-lota-20 for more details.
@@
There is a set $\Pi^{+}$ of $\mathcal{O}(\mathcal{E}^{d}\log\mathcal{E})$ orderings of $[0,1)^{d}$, such that for any two points $p,q\in[0,1)^{d}$ there is an ordering $\sigma\in\Pi^{+}$ defined over $[0,1)^{d}$, such that for any point $u$, with $p\prec_{\sigma}u\prec_{\sigma}q$, we have that either $\left\|{p-u}\right\|\leq\varepsilon\left\|{p-q}\right\|$ or $\left\|{q-u}\right\|\leq\varepsilon\left\|{p-q}\right\|$.
@@ 2.3 Coverings and packings
Consider a set $P\subseteq\mathbb{R}^{d}$, and a parameter $r$. A set $N\subseteq P$ is an $r$-covering of $P$, if for all $p\in P$, there exists a point $p^{\prime}\in N$, such that $\left\|{p-p^{\prime}}\right\|\leq r$. A point set $N$ is $r$-separated if for any two distinct points $p,q\in N$, we have $\left\|{p-q}\right\|>r$. A subset $N\subseteq P$ that is both $r$-separated and an $r$-covering of $P$ is an $r$-packing [^1] Confusingly $r$-packings are also known as $r$-nets, but as the $r$-net is an overloaded concept, we use the alternative, hopefully less-confusing, term. of $P$.
For a finite set $P$ in $\mathbb{R}^{d}$, an $r$-packing can be computed in linear time hr-nplta-15 (for constant $d$). We need such a packing of the unit sphere.
@@
The set $N^{\prime}$ has (asymptotically) the right size. Converting it into an $r$-packing can be done in linear time using the algorithm mentioned above hr-nplta-15.
@@ 3.1 Projections
Let $R>\tau>0$ be parameters. One can compute, in $O(\left|{\mathcal{D}}\right|)=O\bigl{(}(R/\tau)^{d-1}\bigr{)}$ time, a set $\mathcal{D}$ of directions. We have that, for all $p,q\in\mathbb{R}^{d}$, with $\left\|{p-q}\right\|\leq R$, there exists a direction $v\in\mathcal{D}$, such that $\mathsf{d}_{v}\left({p,q}\right)\leq\tau$.
@@
Let $\mathcal{D}$ be an $\xi$-packing of the unit sphere $\mathbb{S}$ computed by Lemma 2.10, where $\xi=\min(\tau/R,1/4)$. We have $\left|{\mathcal{D}}\right|=O(1/\xi^{d-1})=O\bigl{(}(R/\tau)^{d-1}\bigr{)}$. As for correctness, consider the line $\mathcalb{l}$ connecting $p$ to $q$. Let $u$ be the direction vector of $\mathcalb{l}$, and let $v\in\mathcal{D}$ be the closest direction to $u$. By construction, $\left\|{u-v}\right\|\leq\xi$. Let $\mathcalb{g}$ be the line passing through $p$ in the direction of $v$, and let $q^{\prime}$ be the projection of $q$ to this line, see Figure 3.1.
Figure 3.1: Illustration of the proof of Lemma 3.2.
Clearly, $\mathsf{d}_{v}\left({p,q}\right)=\left\|{q-q^{\prime}}\right\|=\left\|{p-q}\right\|\sin\angle qpq^{\prime}\leq R\sin\xi\leq R\xi\leq R\frac{\tau}{R}=\tau,$ since $\left\|{p-q}\right\|\leq R$, $\sin x\leq x$ for $x$ positive, $\xi>0$, and $\xi\leq\tau/R$.
@@ 3.2 Constructing the ordering
    $\left\|{\mathsf{c}\left({\square_{3}}\right)-\mathsf{c}\left({\square_{4}}\right)}\right\|\leq\left\|{\mathsf{c}\left({\square_{1}}\right)-\mathsf{c}\left({\square_{2}}\right)}\right\|$,
@@
Here $\mathsf{c}\left({\square}\right)$ denotes the center point of $\square$.
@@
We now describe this in more detail. Fixing the values of $v$ and $i$, identifies a set $N_{v}^{i}$. For $j=1,\ldots,\left|{N_{v}^{i}}\right|$, let $p_{j}$ be a point of $N_{v}^{i}$ not handled yet. The algorithm computes the oriented line $\mathcalb{l}_{j}$ passing through $p_{j}$ in the direction of $v$. Let $Z_{j}$ be the set of all the cells of $\mathcal{G}$ that $\mathcalb{l}_{j}$ stabs, and their center is in distance at most $\tau$ from $\mathcalb{l}_{j}$. All such cells intersect $\mathcalb{l}$ in a “long” interval of length $\geq\zeta-2\tau\geq(3/4)\zeta\geq 6\tau$ (that is, all the cells of $Z_{j}$ have their centers “almost” on $\mathcalb{l}_{j}$). Thus, the distance between the centers of the cells of $Z_{j}$ in the original space, or the distance between their corresponding projections on $\mathcalb{l}_{j}$ are the same up to ±$\pm 2\tau$. In particular, we sort the cells of $Z_{j}$ according to the order of their projected centers along $\mathcalb{l}_{j}$, and append them (in this sorted order) to the computed order. This process is illustrated in Figure 3.2. Note that in the end of this process, most grid cells of $\mathcal{G}$ are not included in the computed order. We append all these yet unordered cells in an arbitrary fashion to the end of the ordering. Let $\sigma(v,i)$ denote the resulting ordering of the cells of $\mathcal{G}$.
@@ 3.3 An ordering with a gap
and $N=\left|{\mathcal{G}_{T}}\right|=\Theta\bigl{(}(t/\beta)^{d})=\Theta\bigl{(}(t/\alpha)^{d}).$ We partition every cell $\square$ of $\mathcal{G}_{T}$ into a secondary grid $\mathcal{G}_{\square}=\mathcal{G}(\square,\beta)$, see Figure 3.3. After applying this to all the grid cells of $\mathcal{G}_{T}$, this results in the original grid $\mathcal{G}$. Let $\mathfrak{O}_{T}$ be the set of orderings, of Lemma 2.6, of $\mathcal{G}_{T}$ that has the property that every two cells of $\mathcal{G}_{T}$ are adjacent in one of the orderings, where $\left|{\mathfrak{O}_{T}}\right|=O(N)$. Let $\mathfrak{O}_{B}$ be the set of orderings for the original grid $\mathcal{G}$ from Lemma 3.4. Consider an order $\sigma_{T}\in\mathcal{G}_{T}$ and an order $\sigma_{B}\in\mathfrak{O}_{B}$. They induce an ordering of the cells of $\mathcal{G}$ as follows – the top ordering $\sigma_{T}$ orders the cells of $\mathcal{G}$ into secondary blocks, where each secondary block is all the cells of $\mathcal{G}$ that lie in a single cell of $\mathcal{G}_{T}$. Every such block is then sorted using the order $\sigma_{B}$. This defines a new set of orders $\mathfrak{O}$ over $\mathcal{G}$ of size $O(\left|{\mathfrak{O}_{T}}\right|\cdot\left|{\mathfrak{O}_{B}}\right|)=O(t^{2d-1}/\alpha^{d})$.
@@ 3.4 Result
Let $\varepsilon,\gamma\in(0,1/2]$ be fixed constants, such that $\gamma\geq\varepsilon$, and let $\mathcal{E}=1/\varepsilon$. Then, there is a set $\Pi$ of $m=O\bigl{(}(\mathcal{E}^{d-1}\log\mathcal{E})/\gamma^{d}\bigr{)}$ orderings, of $[0,1)^{d}$,such that for any two points $p,q\in[0,1)^{d}$, there is an ordering $\sigma\in\Pi$ defined over $[0,1)^{d}$, such that
@@
    for all points $u$, with $p\prec_{\sigma}u\prec_{\sigma}q$, we have $\mathsf{d}\left({u,pq}\right)\leq\varepsilon\left\|{p-q}\right\|$ and $\mathsf{d}\left({u,\{p,q\}}\right)\leq\gamma\left\|{p-q}\right\|$,
@@
    for all points $u,v$ with $p\prec_{\sigma}u\prec_{\sigma}v\prec_{\sigma}q$, we have $\left\|{u-v}\right\|\leq(1+\varepsilon)\left\|{p-q}\right\|$.
@@
$\mathcal{E}=1/\varepsilon^{\prime},\qquad\text{and}\qquad\alpha=\max\bigl{(}1,\left\lfloor{\gamma\mathcal{E}/(2d^{2})}\right\rfloor\bigr{)}$
Let $\mathfrak{O}=\mathfrak{O}\left({\mathcal{E},\alpha,d}\right)$ be the set of orderings of $\mathcal{G}(\mathcal{C},\mathcal{E})$ as defined by Lemma 3.5. As described in Section 2.2.1, the set $\mathfrak{O}$ induces a set $\Pi=\Pi(\varepsilon^{\prime},\alpha)$ of orderings of $[0,1)^{d}$, where
@@
So consider any two points $p,q\in[0,1)^{d}$. There is a shift $\nu\in\mathcal{S}$, and an $\varepsilon^{\prime}$-quadtree $\mathcal{T}_{\varepsilon}\in\mathcal{F}$, such that the LCA of $p$ and $q$ in $\mathcal{T}_{\varepsilon}$ is a node $w$, such that $\mathrm{sidelen}\left({\square_{w}}\right)\leq(d+1)\left\|{p-q}\right\|$, see Lemma 2.5, where $\mathrm{sidelen}\left({\square}\right)=\mathrm{diam}\left({\square}\right)/\sqrt{d}$ is the sidelength of $\square$. In particular, let $\zeta=\mathrm{sidelen}\left({\square_{w}}\right)/\mathcal{E}$ be the sidelength of the subcells of the grid of $\mathcal{T}_{\varepsilon}$ for $\square_{w}$. Observe that $\left\|{p-q}\right\|\geq\mathcal{E}\zeta/(d+1)$. Let $\square_{1}$ and $\square_{2}$ be the cells of the two children of $w$ that contains $p$ and $q$, respectively. In particular, we have that
@@
Computing the orderings of Theorem 3.7 can be done in $\mathcal{E}^{O(d)}$ time, and since this is a preprocessing stage, we ignore this running time in the theorem statement. This results in $m=O\bigl{(}(\mathcal{E}^{d-1}\log\mathcal{E})/\gamma^{d}\bigr{)}$ orderings, each order requires $O(\mathcal{E}^{d})$ space to store it. In the following, $\gamma$ is a fixed “large” constant (e.g., $\gamma=1/8$), so the overall space to store these orderings is $O(\mathcal{E}^{2d-1}\log\mathcal{E})$. Given an order $\sigma$ and two points, comparing the two points according to $\sigma$ requires answering a single LCA query on a quadtree, see Section 2.2.1. With the appropriate bit operations such a query can be carried out in $O(\log\mathcal{E})$ time chj-lota-20.